# An Analytical Cost Model for Fast Evaluation of Multiple Compute-Engine CNN Accelerators

### v2

@@ Abstract
This paper proposes a Multiple-CE accelerator analytical Cost Model (MCCM) and an evaluation methodology built around MCCM. The model and methodology streamline the expression of any multiple-CE accelerator and provide a fast evaluation of its performance and efficiency. MCCM is in the order of 100000×100000$\times$100000 × faster than traditional synthesis-based evaluation and has an average accuracy of $>90\%$. The paper presents three use cases of MCCM. The first describes an end-to-end evaluation of state-of-the-art multiple-CE accelerators considering various metrics, CNN models, and resource budgets. The second describes fine-grained evaluation that helps identify performance bottlenecks of multiple-CE accelerators. The third demonstrates that MCCM fast evaluation enables exploring the vast design space of multiple-CE accelerators. These use cases show that no unique CE arrangement achieves the best results given different metrics, CNN models, and resource budgets. They also show that fast evaluation enables design space exploration, resulting in accelerator designs that outperform state-of-the-art ones. MCCM is available at https://github.com/fqararyah/MCCM
@@ I Introduction
A systematic accelerator design approach requires analyzing the bottlenecks and inefficiencies and exploring the design space to identify optimizations that mitigate them. The literature reveals that the inefficiencies of Deep Learning (DL) accelerators, in general, have three root causes, (1) Processing Element (PE) underutilization [4, 23, 48], (2) large on-chip buffers [4, 27, 46], (3) and the time and energy costly off-chip access [48, 23]. Considering multiple-CE accelerators, the second and third causes are still open challenges. However, multiple-CE accelerators have less PE underutilization than generic ones. Multiple-CE accelerators mitigate PE underutilization bottleneck, but in a way that creates throughput and latency trade-off [41, 30, 5]. Hence, a systematic multiple-CE accelerator design approach requires analyzing PE underutilization with latency-throughput trade-offs, on-chip buffer requirements, and off-chip access.
@@
MCCM is in the order of 100000×100000$\times$100000 × faster than synthesis-based evaluation and has an average accuracy of $>90\%$. The presented MCCM use cases using three state-of-the-art multiple-CE accelerators, five CNNs, and four FPGA boards show that no single multiple-CE architecture is the best given different metrics, CNN models, and resource budgets. Moreover, the evaluation shows the impact of engine arrangements on multiple-CE accelerators’ performance bottlenecks. Finally, using MCCM fast evaluation as a basis for multiple-CE accelerators design space exploration enables identifying designs that outperform the state-of-the-art.
@@ II-A Convolutional Neural Networks (CNNs)
Convolutional Neural Networks (CNNs) are feed-forward Deep learning (DL) algorithms [20]. A CNN comprises a sequence of layers, that perform feature extraction and classification [21]. Convolutional layers are the primary layers in a CNN. A convolutional layer consists of a collection of trainable parameters, known as weights, which are organized into filters. These filters extract features from multi-dimensional inputs or intermediate results. The inputs of a layer are referred to as Input Feature Maps (IFMs), and the outputs as Output Feature Maps (OFMs). The term Feature Maps (FMs), or activations, refers to both IFMs and OFMs. FMs consist of 2D slices known as channels.
@@ II-B CNN Compute-Engines
As convolutions represent more than 90% of modern CNN operations [23], CNN Compute-Engine (CE) optimizations focus on convolutional layers. A CE comprises a grid of Processing Elements (PEs) where each PE performs a MAC operation, e.g. a DSP in an FPGA. Two main design decisions that affect the performance and efficiency of a CE are parallelism strategy and dataflow. A convolution operation comprises a loop nest of six loops, without batching [17]. A CE parallelism strategy describes which among these six loops are partially or fully parallelized. The CE parallelism strategy affects data reuse and memory access patterns [8, 23]. Figure 1 depicts examples of parallelism strategies in 3-D, 2-D, and 1-D. An exhaustive analysis of parallelism strategies on FPGAs shows that parallelizing on three dimensions, namely across filters and within an IFM channel width and height, yields the best average reuse and CE utilization across the layers of common CNNs [23]. However, 2-D and 1-D parallelism are also common when each CE has a limited PEs budget, or when 1-D or 2-D options are more compatible with the dimensions of the CNN layers processed by that CE [47, 24, 1, 33, 49, 37, 50]. In the context of convolution CEs, the term dataflow refers to both convolution loop transformations, i.e. the scheduling of the loops computations, and the mapping of computations across the CE PEs. The most common dataflow types are weight-stationary, output-stationary, and input-stationary. Each dataflow indicates which of the IFMs, OFMs, or weights is scheduled to move least frequently [18].
@@ II-C Multiple compute-engine CNN accelerators
Figure 2: CNN to multiple-CE architecture mapping examples. For example, in Segmented, CE1 processes layers L1-L4, CE2 processes layers L5 and L6, and so on. In practice, CE and buffer sizes are proportional to the segment layers’ compute and memory requirements respectively
@@
We refer to the first architecture in Figure 2 as Segmented [32, 33]. In this architecture, each CE processes one or more segments. Each segment consists of a set of consecutive layers. Each CE has its on-chip buffer that stores the weights and FMs partially or fully depending on the available on-chip memory. Coarse-grained pipelining, where different CEs process different inputs (images) at a certain time step, is applied between the CEs. CE inputs are double-buffered to enable pipelining. We call the second architecture SegmentedRR (round-robin) as CEs in such architecture usually process CNN layers circularly according to their topological order [38, 3, 41]. While some apply course-grained pipelining between CEs, Wei et. al [41] have shown that, for this architecture, tile-grained pipelining is more efficient. In addition to the double-buffering between the CEs, each CE has a separate buffer for the weights. We call the third architecture Hybrid [1, 50, 25, 30]. It comprises two parts. The first contains a set of tile-grained pipelined CEs each processing a single layer. The second has a larger CE that processes the rest of the layers. If the CNN has two types of convolutional layers the second part could have two sub-CEs [30]. Coarse-grained pipelining is applied between Hybrid’s first and second parts.
@@ II-D Multiple-CE accelerator modeling and evaluation
State-of-the-art multiple-CE accelerators adopt one variant of the three mentioned architectures and focus on optimizing over two extremes, namely reusable-CEs based accelerators [2, 23, 45] and accelerators with as many CEs as CNN layers [49, 36]. Adopting fixed architectures limits the potential of existing multiple-CE accelerators. No single architecture is optimal when considering different metrics. Table I, which reports normalized synthesis results of instances of the three architectures, exemplifies that. For example, the SegmentedRR at the first row has the best latency but requires 2.64×2.64$\times$2.64 × the on-chip buffer and 1.79×1.79$\times$1.79 × the off-chip accesses of the best accelerator in each metric. Note that buffer requirements and off-chip memory accesses are treated here as optimization goals. This is because minimizing off-chip accesses, consequently bandwidth requirements, and on-chip buffer requirements are crucial, especially when resources are shared by multiple applications [40], even when neither of them forms a bottleneck. The fact that no single architecture gives the best results considering different metrics, even for the same CNN and FPGA as the table shows, suggests that a fixed architecture is far from optimal. Hence, a more systematic design approach that explores different CE arrangement possibilities is needed to identify the best designs given an application performance metrics, CNN model, and resource budget. However, a traditional evaluation of one CE arrangement can take hours. This makes exploring the numerous CE arrangement possibilities of multiple-CE accelerators, and consequently a systematic design approach, impractical. This motivates proposing a fast evaluation methodology of multiple-CE accelerators.
@@ III-A Overview of the evaluation methodology
Figure 3 presents an overview of the proposed multiple-CE accelerator evaluation methodology. The methodology takes as inputs: (1) multiple-CE accelerator description, discussed in the Section III-B, (2) a CNN representation, (3) FPGA’s number of PEs (DSPs), off-chip memory bandwidth, and on-chip memory capacity. The main outputs are throughput, latency, on-chip buffer requirements, and off-chip accesses. Other outputs include a fine-grained analysis of PE utilization and a breakdown of the results on the level of weights and FMs. The evaluation methodology comprises three modules. The Multiple-CE Blocks module contains codified descriptions of two building blocks that could be used to construct any multiple-CE accelerator, namely single-CE and pipelined-CEs. A building block comprises one or more CEs. The Multiple-CE Builder module constructs a generic multiple-CE accelerator backbone by combining the building blocks considering the three methodology inputs. It then decides the implementation details, including the CEs buffer and PE distribution, dataflows, and parallelism strategies based on a set of heuristics inspired by the prior art [33, 3, 41, 23, 30]. The MCCM module evaluates a generic multiple-CE accelerator bottom-up by modeling its building blocks and their interfaces. MCCM estimates multiple-CE accelerator throughput, latency, on-chip buffers, and memory access requirements. MCCM is the main contribution of this work; hence, we use the terms MCCM and evaluation methodology interchangeably. As DL accelerator design is a hot topic where new accelerators are frequently proposed, we designed MCCM to be modular and extensible. MCCM interface allows easy addition of models of newly proposed specialized designs. MCCM components are discussed in detail throughout Section IV.
@@ IV-A3 Off-chip access
$\begin{split}Accesses&=\sum_{i=1}^{layers}{Acc(L_{i},\textit{CE}_{j})}\\ \textit{Acc(L}_{i},\textit{CE}_{j})&=argmin\bigg{(}\textit{OFMsSz}_{i}\times\textit{offCh(OFMs}_{i})+\\ &\textit{min}\Bigl{(}\textit{weightsSz}_{i}\ \times\frac{\textit{IFMsSz}_{i}}{\textit{IFMsBufferSz}_{j}}+\\ &\textit{IFMsSz}_{i},\textit{IFMsSz}_{i}\ \times\frac{\textit{weightsSz}_{i}}{\textit{weightsBufferSz}_{j}}+\\ &\textit{weightsSz}_{i}\Bigl{)}\times\textit{offCh(IFMs}_{i})+\\ &\Bigl{(}1-\textit{offCh(IFMs}_{i})\Bigl{)}\times\textit{weightsSz}_{i}\bigg{)}\\ s.t.\ \textit{weights}&\textit{BufferSz}_{j}+\textit{IFMsBufferSz}_{j}+\max_{i=1}^{layers}\bigl{(}\textit{OFMsSz}_{i}\times\\ &(1-\textit{offCh(OFMs}_{i}))\bigl{)}\leq\textit{CE\_BufferSz}_{j}\end{split}$ (6)
@@ V-A2 Workloads
We use a representative set of CNNs with different structures, depths, and parameter (weight) counts. These are ResNet152 and ResNet50 [15], XCeption  [10], MobileNetV2 [31], and DenseNet [16]. Table III shows the relevant characteristics of these CNNs. The evaluated CNNs share core blocks with many recent CNNs. For example, MBConv, the core block of MobileNetV2, is used in EfficientNet [35] and MnasNet [34]. The same applies to the Residual blocks of ResNets, and the Extreme Inceptions of XCeption are also used in recent CNNs.

### v3

@@ Abstract
This paper proposes a Multiple-CE accelerator analytical Cost Model (MCCM) and an evaluation methodology built around MCCM. The model and methodology streamline the expression of any multiple-CE accelerator and provide a fast evaluation of its performance and efficiency. MCCM is in the order of 100000×100000$\times$100000 × faster than traditional synthesis-based evaluation and has an average accuracy of $>90\%$. The paper presents three use cases of MCCM. The first describes an end-to-end evaluation of state-of-the-art multiple-CE accelerators considering various metrics, CNN models, and resource budgets. The second describes fine-grained evaluation that helps identify performance bottlenecks of multiple-CE accelerators. The third demonstrates that MCCM fast evaluation enables exploring the vast design space of multiple-CE accelerators. These use cases show that no unique CE arrangement achieves the best results given different metrics, CNN models, and resource budgets. They also show that fast evaluation enables design space exploration, resulting in accelerator designs that outperform state-of-the-art ones. MCCM is available at https://github.com/fqararyah/MCCM.
@@ I Introduction
A systematic accelerator design approach requires analyzing the bottlenecks and inefficiencies and exploring the design space to identify optimizations that mitigate them. The literature reveals that the inefficiencies of Deep Learning (DL) accelerators, in general, have three root causes, (1) Processing Element (PE) underutilization [4, 23, 48], (2) large on-chip buffers [4, 27, 46], (3) and the time and energy costly off-chip access [48, 23]. Considering multiple-CE accelerators, the second and third causes are still open challenges. However, multiple-CE accelerators have less PE underutilization than generic ones. Multiple-CE accelerators mitigate PE underutilization bottleneck but in a way that creates throughput and latency trade-off [41, 30, 5]. Hence, a systematic multiple-CE accelerator design approach requires analyzing PE underutilization with latency-throughput trade-offs, on-chip buffer requirements, and off-chip access.
@@
MCCM is in the order of 100000×100000$\times$100000 × faster than synthesis-based evaluation and has an average accuracy of $>90\%$. The presented MCCM use cases using three state-of-the-art multiple-CE accelerators, five CNNs, and four FPGA boards show that no single multiple-CE architecture is the best, given different metrics, CNN models, and resource budgets. Moreover, the evaluation shows the impact of engine arrangements on multiple-CE accelerators’ performance bottlenecks. Finally, using MCCM fast evaluation as a basis for multiple-CE accelerators design space exploration enables identifying designs that outperform the state-of-the-art.
@@ II-A Convolutional Neural Networks (CNNs)
Convolutional Neural Networks (CNNs) are feed-forward Deep learning (DL) algorithms [20]. A CNN comprises a sequence of layers that perform feature extraction and classification [21]. Convolutional layers are the primary layers in a CNN. A convolutional layer consists of a collection of trainable parameters, known as weights, which are organized into filters. These filters extract features from multi-dimensional inputs or intermediate results. The inputs of a layer are referred to as Input Feature Maps (IFMs), and the outputs as Output Feature Maps (OFMs). The term Feature Maps (FMs), or activations, refers to both IFMs and OFMs. FMs consist of 2D slices known as channels.
@@ II-B CNN Compute-Engines
As convolutions represent more than 90% of modern CNN operations [23], CNN Compute-Engine (CE) optimizations focus on convolutional layers. A CE comprises a grid of Processing Elements (PEs) where each PE performs a MAC operation, e.g., a DSP in an FPGA. Two main design decisions that affect the performance and efficiency of a CE are parallelism strategy and dataflow. A convolution operation comprises a loop nest of six loops without batching [17]. A CE parallelism strategy describes which among these six loops are partially or fully parallelized. The CE parallelism strategy affects data reuse and memory access patterns [8, 23]. Figure 1 depicts examples of parallelism strategies in 3-D, 2-D, and 1-D. An exhaustive analysis of parallelism strategies on FPGAs shows that parallelizing on three dimensions, namely across filters and within an IFM channel width and height, yields the best average reuse and CE utilization across the layers of common CNNs [23]. However, 2-D and 1-D parallelism are also common when each CE has a limited PEs budget or when 1-D or 2-D options are more compatible with the dimensions of the CNN layers processed by that CE [47, 24, 1, 33, 49, 37, 50]. In the context of convolution CEs, the term dataflow refers to both convolution loop transformations, i.e. the scheduling of the loop computations, and the mapping of computations across the CE PEs. The most common dataflow types are weight-stationary, output-stationary, and input-stationary. Each dataflow indicates which of the IFMs, OFMs, or weights is scheduled to move least frequently [18].
@@ II-C Multiple compute-engine CNN accelerators
Figure 2: CNN to multiple-CE architecture mapping examples. For example, in Segmented, CE1 processes layers L1-L4, CE2 processes layers L5 and L6, and so on. In practice, CE and buffer sizes are proportional to the segment layers’ compute and memory requirements, respectively.
@@
We refer to the first architecture in Figure 2 as Segmented [32, 33]. In this architecture, each CE processes one or more segments. Each segment consists of a set of consecutive layers. Each CE has its on-chip buffer that stores the weights and FMs partially or fully, depending on the available on-chip memory. Coarse-grained pipelining, where different CEs process different inputs (images) at a certain time step, is applied between the CEs. CE inputs are double-buffered to enable pipelining. We call the second architecture SegmentedRR (round-robin) as CEs in such architecture usually process CNN layers circularly according to their topological order [38, 3, 41]. While some apply course-grained pipelining between CEs, Wei et. al [41] have shown that, for this architecture, tile-grained pipelining is more efficient. In addition to the double-buffering between the CEs, each CE has a separate buffer for the weights. We call the third architecture Hybrid [1, 50, 25, 30]. It comprises two parts. The first contains a set of tile-grained pipelined CEs, each processing a single layer. The second has a larger CE that processes the rest of the layers. If CNN has two types of convolutional layers, the second part could have two sub-CEs [30]. Coarse-grained pipelining is applied between the Hybrid’s first and second parts.
@@ II-D Multiple-CE accelerator modeling and evaluation
State-of-the-art multiple-CE accelerators adopt one variant of the three mentioned architectures and focus on optimizing over two extremes, namely reusable-CEs based accelerators [2, 23, 45] and accelerators with as many CEs as CNN layers [49, 36]. Adopting fixed architectures limits the potential of existing multiple-CE accelerators. No single architecture is optimal when considering different metrics. Table I, which reports normalized synthesis results of instances of the three architectures, exemplifies that. For example, the SegmentedRR at the first row has the best latency but requires 2.64×2.64$\times$2.64 × the on-chip buffer and 1.79×1.79$\times$1.79 × the off-chip accesses of the best accelerator in each metric. Note that buffer requirements and off-chip memory accesses are treated here as optimization goals. This is because minimizing off-chip accesses, consequently bandwidth requirements, and on-chip buffer requirements are crucial, especially when resources are shared by multiple applications [40], even when neither of them forms a bottleneck. The fact that no single architecture gives the best results considering different metrics, even for the same CNN and FPGA as the table shows, suggests that a fixed architecture is far from optimal. Hence, a more systematic design approach that explores different CE arrangement possibilities is needed to identify the best designs given an application performance metrics, CNN model, and resource budget. However, a traditional evaluation of one CE arrangement can take hours. This makes exploring the numerous CE arrangement possibilities of multiple-CE accelerators, and consequently a systematic design approach, impractical. This motivates the proposal of a fast evaluation methodology for multiple-CE accelerators.
@@ III-A Overview of the evaluation methodology
Figure 3 presents an overview of the proposed multiple-CE accelerator evaluation methodology. The methodology takes as inputs: (1) multiple-CE accelerator description, discussed in the Section III-B, (2) a CNN representation, (3) FPGA’s number of PEs (DSPs), off-chip memory bandwidth, and on-chip memory capacity. The main outputs are throughput, latency, on-chip buffer requirements, and off-chip accesses. Other outputs include a fine-grained analysis of PE utilization and a breakdown of the results on the level of weights and FMs. The evaluation methodology comprises three modules. The Multiple-CE Blocks module contains codified descriptions of two building blocks that could be used to construct any multiple-CE accelerator, namely single-CE and pipelined-CEs. A building block comprises one or more CEs. The Multiple-CE Builder module constructs a generic multiple-CE accelerator backbone by combining the building blocks considering the three methodology inputs. It then decides the implementation details, including the CEs buffer and PE distribution, dataflows, and parallelism strategies based on a set of heuristics inspired by the prior art [33, 3, 41, 23, 30]. The MCCM module evaluates a generic multiple-CE accelerator bottom-up by modeling its building blocks and their interfaces. MCCM estimates multiple-CE accelerator throughput, latency, on-chip buffers, and memory access requirements. MCCM is the main contribution of this work; hence, we use the terms MCCM and evaluation methodology interchangeably. MCCM components are discussed in detail throughout Section IV. As DL accelerator design is a hot topic where new accelerators are frequently proposed, we designed MCCM to be modular and extensible. MCCM interface allows easy addition of models of newly proposed specialized designs. [^1] The details on how to extend MCCM are provided in the repository: https://github.com/fqararyah/MCCM.
@@ IV-A3 Off-chip access
$\begin{split}Accesses&=\sum_{i=1}^{layers}{Acc(L_{i},\textit{CE}_{j})}\\ \textit{Acc(L}_{i},\textit{CE}_{j})&=argmin\bigg{(}\textit{OFMsSz}_{i}\times\textit{offCh(OFMs}_{i})+\\ &\textit{min}\Bigl{(}\textit{weightsSz}_{i}\ \times\lceil\frac{\textit{IFMsSz}_{i}}{\textit{IFMsBufferSz}_{j}}\rceil+\\ &\textit{IFMsSz}_{i},\textit{IFMsSz}_{i}\ \times\lceil\frac{\textit{weightsSz}_{i}}{\textit{weightsBufferSz}_{j}}\rceil+\\ &\textit{weightsSz}_{i}\Bigl{)}\times\textit{offCh(IFMs}_{i})+\\ &\Bigl{(}1-\textit{offCh(IFMs}_{i})\Bigl{)}\times\textit{weightsSz}_{i}\bigg{)}\\ s.t.\ \textit{weights}&\textit{BufferSz}_{j}+\textit{IFMsBufferSz}_{j}+\max_{i=1}^{layers}\bigl{(}\textit{OFMsSz}_{i}\times\\ &(1-\textit{offCh(OFMs}_{i}))\bigl{)}\leq\textit{CE\_BufferSz}_{j}\end{split}$ (6)
@@ V-A2 Workloads
We use a representative set of CNNs with different structures, depths, and parameter (weight) counts. These are ResNet152 and ResNet50 [15], XCeption  [10], MobileNetV2 [31], and DenseNet [16]. Table III shows the relevant characteristics of these CNNs. The results presented can be generalized as the evaluated CNNs share core blocks with many recent CNNs. For example, MBConv, the core block of MobileNetV2, is used in EfficientNet [35] and MnasNet [34]. The same applies to the Residual blocks of ResNets, and the Extreme Inceptions of XCeption are also used in recent CNNs.
@@ VII Conclusion
Acknowledgement
This work was supported by the VEDLIoT project, which received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 957197. This work was also partly supported by the Swedish Foundation for Strategic Research (contract number CHI19-0048) under the PRIDE project and the European High-Performance Computing Joint Undertaking (JU) under Framework Partnership Agreement No 800928 and Specific Grant Agreement No 101036168 (EPI SGA2). The JU receives support from the European Union’s Horizon 2020 research and innovation program and from Croatia, France, Germany, Greece, Italy, Netherlands, Portugal, Spain, Sweden, and Switzerland.